the measure of angle j is half of its compliment plus 12 degrees: the measure of angle j is half of its compliment plus 12 degrees
Answer by mangopeeler07(442)

j=1/2(c+12)
c=c
j+c=90
1/2(c+12)+c=90

1/2c+6+c=90
3/2c+6=90

3/2c=84
c=55.99
j=90-c
j=90-55.99
j=34.01

j=34.01
j's compliment=55.99

------------------------------------OR----------------------------------------
j=(1/2c)+12
c=c

j+c=90
1/2c+12+c=90

3/2c+12=90
3/2c=78
c=52

j=38
j's compliment=52
Question 149389: the measure of angle j is half of its compliment plus 12 degrees: the measure of angle j is half of its compliment plus 12 degrees
Answer by jim_thompson5910(9165)

"measure of angle j is half of its compliment plus 12 degrees" translates to j=(1/2)(c+12)


Also, since angle j and angle c are complementary angles, this means that j+c=90


(1/2)(c+12)+c=90 Plug in j=(1/2)(c+12)


c+12+2c=180 Multiply every term by 2 to clear the fraction.


c+2c=168 Subtract 12 from both sides.


3c=168 Combine like terms.


c=56 Divide both sides by 3.


j=(1/2)(c+12) Go back to the first equation.


j=(1/2)(56+12) Plug in c=56


j=(1/2)(68) Add


j=34 Multiply.